OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (2, -9) and (4, 1). Please show all of your work.

OpenStudy (anonymous):

Find the slope first. Then use one point with slope to make the line. Do you know how to find slope

OpenStudy (anonymous):

\[y - y _{1} = m (x - x _{1})\] \[m = y _{2} - y _{1} \div x _{2} - x _{1}\]